The post-synaptic cell at a chemical synapse has the following distribution of ions between the cytosol and the extracellular***I need help on PART E of this problem. All of the information is given on the worksheet along with the hints on each part. Thank you!
Answer #1

Abswer a)

a) When ions of opposite charge are separated by á permeable membrane, there is a transmembrane electrochemical gradient, a m

AR 25°C, V 500 96480 = 3 3 15 x 218K .( 10p(5) . (146) 96480mn 1 100 (150) = 0.025 in (0.033 + 29.2 + 0.032) = 0.025 in (29.2

Answer b)

til it exactly b) Extra cellular and Intracellular concentrations of ions are given - charge separation across the membrane l

for K, V = 0.025 in -0.085 =[- 85 m - For I, = 2477.87 155 0.025 TD 96480 - - 0.025 in (31) - 0.085 = 85ml

Since the membrane potential of indiviaual ions are either positive or negative (and not zero), none of the ions are at electrochemical equilibrium.

Answer c)

Pk = 200 c) Permeability of potassium changes to Using GHK ean - 1200 (5) + 29.2+ Vm = 0.025 en 200 (150) = 3.376 v OR 3370 mSo, changing the ion permeability does not change membrane potential.

Answer d)

d) Nm = 0.025 in (0.033 + 29.2 + 20015)) = 3.376 V OR 3376 mVSo, changing the ion permeability does not change membrane potential.